radical sign over 4x-3 - raddical sign over 2x+7 =0 solve equation
Answer by ankor@dixie-net.com(22740)

radical sign over 4x-3 - radical sign over 2x+7 = 0 solve equation
:
SqRt(4x-3) - SqRt(2x+7) = 0
:
Add SqRt(2x+7) to both sides and you have:
SqRt(4x-3) = SqRt(2x+7)
:
Square both sides:
4x - 3 = 2x + 7
:
4x - 2x = 7 + 3
:
2x = 10
:
x = 5
:
Check: SqRt(20 - 3) - SqRt(10 + 7) = 0